OpenStudy (mayankdevnani):

A triangle is formed by the lines x+y=0 , x-y=0 and lx+my=1 . If l and m are the parameters such that \(\large \bf l^2+m^2=1\) , then the locus of its circumcenter is

OpenStudy (mayankdevnani):

Options are :- \[\large \bf A) (x^2-y^2)^2=x^2+y^2\] \[\large \bf B) (x^2+y^2)^2=(x^2-y^2)\] \[\large \bf C) (x^2+y^2)^2=4x^2y^2+(x^2+y^2)\] \[\large \bf D) (x^2-y^2)^2=(x^2+y^2)^2\]

hartnn (hartnn):

this doesn't seem difficult, what have you tried?

hartnn (hartnn):

 x,y is just the mid-point of 2 intersection points, I1, I2.